I am suppose to solve the following system but don't know how.
-2x+2y+3z=-7 (1)
x-2y-2z=4 (2)

Answer by 303795(602)

-2x+2y+3z=-7 (1)
x-2y-2z=4 (2)
You have three unknowns but only two equations. That is not enough to solve for x, y and z.
You can add the two equations to get -x + z = -3 but that is as far as you can go without another equation.